1. 1. Find a basis for the subspace of R Superscript 3 that is spanned by the vectors

v1 = (1, 0, 0), v2 = (1, 1, 0), v3 = (3, 1, 0), v4 = (0, -2, 0). Choose the correct letter.

A. v1 and v2 form a basis for span {v1, v2, v3, v4}.
B . v2 and v3 form a basis for span {v1, v2, v3, v4}.
C. v3 and v4 form a basis for span {v1, v2, v3, v4}.
D. v1 and v3 form a basis for span {v1, v2, v3, v4}.
E. v2 and v4 form a basis for span {v1, v2, v3, v4}.
F. v1 and v4 form a basis for span {v1, v2, v3, v4}.
1
Expert's answer
2020-01-12T14:17:34-0500

"c_1v_1+c_2v_2+c_3v_3+c_4v_4=0"

"c_1(1, 0, 0)+c_2(1, 1, 0)+c_3(3, 1, 0)+c_4(0, -2, 0)=0"

From the above equation we get the following equations:

"c_1+c_2+3c_3=0"

"c_2+c_3-2c_4=0"

Simplifying we get:

"c_1=2c_3-2c_4"

"c_2=2c_4-c_3"

The {"v_1,v_2" } spans span of subspace and are linearly independent.

Indeed,

"c_1v_1+c_2v_2=0," "c_1(1,0,0)+c_2(1,1,0)=0."

We obtain the following equations

"c_1+c_2=0,"

"c_2=0."

Hence "c_2=0, c_1=0."

Thus, {"v_1,v_2" } is the basis.

Checking option (A)

"v_3=2v_1+v_2," "v_4=2v_1-2v_2."

Thus, "v_1" and "v_2" form a basis for span "\{ v_1, v_2, v_3, v_4 \}."